Shift report 20010926 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Wed Sep 26 2001 - 16:05:23 EDT

  • Next message: Flemming Videbaek: "Updated brahms shift list"

    ------------------------------------------------------------
    Date: Wed Sep 26 16:05:22 2001, Shift: 08:00-16:00
    Supervisor: J.I.Jordre, with: none
    Log book#9 pages 49 - 50
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam studies in the morning. Fresh store from 14:34.
    
    2. Run 4968		TPC pedestals
       Run 4969		Pedestal run without TPCs
       Run 4970		DVM run
       Run 4971		Data run, still running
       (MRS @ 90 deg. A pol., FS @ 3 deg. A pol.)
       Dana also did a angular scan for FS.
    
    3. Right ZDC low ADC module 1 gone again. FV just fixed it, but it
       but it is off for run 4971.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Wed Sep 26 2001 - 16:06:00 EDT